# Matching seismic masses for RR Lyrae-type and oscillating red horizontal-branch stars in M4

### v1

@@
11email: molnar.laszlo@csfk.org; lmolnar@konkoly.hu 22institutetext: ELTE Eötvös Loránd University, Institute of Physics and Astronomy, 1117, Pázmány Péter sétány 1/A, Budapest, Hungary 33institutetext: Institute of Physics, École Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, 1290 Versoix, Switzerland 44institutetext: School of Physics and Astronomy, Monash University, Clayton, VIC 3800, Australia 55institutetext: ARC Centre of Excellence for Astrophysics in Three Dimensions (ASTRO-3D), Australia 66institutetext: University of Wyoming, 1000 E University Ave, Laramie, WY USA
@@
Globular clusters offer a powerful way to test the properties of stellar populations and the late stages of low-mass stellar evolution. In this paper we study oscillating giant stars and overtone RR Lyrae-type pulsators in the nearest globular cluster, M4, with the help of high-precision, continuous light curves collected by the Kepler space telescope in the K2 mission. We determine the frequency composition of five RRc stars and model their physical parameters with a grid of linear pulsation models. We are able, for the first time, to compare seismic masses of RR Lyrae stars directly to the masses of the very similar red horizontal branch stars in the same stellar population, independently determined from asteroseismic scaling relations. We find a close match, with an average seismic mass of for RR Lyrae stars and for red horizontal-branch stars. While the validity of our RR Lyrae masses still relies on the similarity of neighboring horizontal branch subgroups, this result strongly indicates that RRc stars may indeed exhibit high-degree, and 9 non-radial modes, and modeling these modes can provide realistic mass estimates. We also determine the He content of the cluster to be , and compare the seismic masses for our sample of RR Lyrae to theoretical mass relations and highlight the limitations of these relations.
@@
Binarity, for example, would be a straightforward way to infer dynamical masses. However, not only do we not know any eclipsing binaries with an RR Lyrae member, it is very difficult to find systems that contain an RR Lyrae star at all (Wade et al., 1999; Kennedy et al., 2014; Hajdu et al., 2015). Since bona fide RR Lyrae stars are considered to be very old, their companions have either evolved into white dwarfs or other stellar remnants already, or must be smaller stars, most likely still evolving along the lower main sequence as K–M dwarfs. As such, the contribution of RR Lyr companions to the total light output of the system is expected to be minimal, precluding the detection via spectral energy distributions or as double-lined spectroscopic binaries.
The dearth of eclipsing RR Lyrae stars can also be explained by the evolutionary pathways of these pulsators. These stars moved past the red giant phase without any significant interactions with a companion that could have directed the primary away from evolving onto the HB and towards binary evolution scenarios. Companions thus must be distant to avoid any mass transfer, and a distant companion orbiting a now shrunken horizontal-branch star would make the probability of observing eclipses from the system very low. The only eclipsing binary candidate system turned out to be an impostor, a Binary Evolution Pulsator (Pietrzyński et al., 2012; Karczmarek et al., 2017). In this case, mass transfer directed a much lighter primary to rapidly cross the instability strip and to experience RR Lyrae-like pulsations. We note that binarity and a limited amount of mass transfer has been proposed to explain metal-rich RR Lyrae stars, but those systems are expected to have wide orbits, too (Bobrick et al., 2024).
More indirect methods have shown increased promise, but all have their own caveats. For example, radial velocity searches are hindered by the expected long orbital periods and small signals relative to the much larger pulsation velocities. A recent work proposed new binary candidates, but it also failed to confirm the expected periastron passage of the oldest binary candidate, TU UMa (Wade et al., 1999; Barnes et al., 2021).
An alternative way to detect binaries and estimate masses is the light-time effect, but this method has been impeded by the prevalence of the Tseraskaya–Blazhko effect111Ms. Tseraskaya (Ceraski) collected the data at Moscow Observatory that led to the discovery., the quasi-periodic modulation of the pulsation amplitude and phase (Blažko, 1907; Shapley, 1916; Kurtz, 2022). Nevertheless, several binary candidates have been identified in the OGLE survey (Hajdu et al., 2015, 2021). And while Hajdu et al. (2021) did not address the masses of the primaries directly, these authors found a trimodal distribution in the mass functions of the secondaries.
@@
One way to limit uncertainties in stellar parameters is to study RR Lyrae stars in globular clusters. There are currently close to three thousand RR Lyrae stars known in 115 Galactic globular clusters (Cruz Reyes et al., 2024). For these stars, parameters like distance and metallicity can be determined accurately for the members, and using that information, the van Albada & Baker (1971) relation was employed in many cases to estimate masses. M3, for example, was studied by multiple authors, and Valcarce & Catelan (2008) determined the mass distribution for the entire HB of the cluster based on evolutionary models, and found the distribution to be potentially bimodal. However, they discuss how that finding can be complicated by limitations present in the stellar tracks. A different approach was taken by Kumar et al. (2024), who fitted the light curves of the M3 RR Lyrae variables directly, based on a grid of non-linear pulsation models. Direct light curve fitting is a promising technique (see also, Bellinger et al., 2020, and Das et al., in prep), but current one-dimensional non-linear pulsation models still cannot reproduce light curves very accurately, due to limitations in handling convection, for example (Kovács et al., 2023).
Uncertainties in evolutionary and non-linear pulsation models can also be mitigated if we turn towards asteroseismology, and model the frequencies of the oscillatory modes observed in the star (see, e.g., Kjeldsen & Bedding, 1995; Aerts, 2021). Incorporating asteroseismic constraints has shown great potential recently for stars outside the instability strip. However, classical pulsators like RR Lyrae stars face obstacles in this aspect, too: in these stars, most often only a single mode can be observed, severely limiting our ability to constrain physical parameters from frequencies alone. Masses for a frew double-mode RR Lyrae stars have been calculated from non-linear non-adiabatic hydrodynamical simulations, (Molnár et al., 2015b), but, as mentioned before, the accuracy of non-linear models have been called into question (Smolec & Moskalik, 2008).
More recently, a new method based on the discovery of low-amplitude signals in overtone RR Lyrae stars has been proposed (Netzel et al., 2015b, c). Theoretical predictions by Dziembowski (2016) suggested that these modes may be high-order, , 9 non-radial modes, which can be calculated from linear pulsation models. Unlike modeling radial modes in the non-linear regime, fitting radial and non-radial modes together using a linear model grid is a technique very similar to the asteroseismology of red giant stars. This method was developed by Netzel & Smolec (2022), and was further explored, incorporating other observables such as , L and [Fe/H] information into the fit in Netzel et al. (2023).
@@
This opportunity came with the Kepler space telescope. Kepler observed multiple globular clusters during the K2 mission, including the cluster closest to the Sun, Messier 4 (M4). This cluster was bright enough that seismic data was obtainable not only for the bright red giants, but fainter stars, such as the rHB stars in it, as well. Howell et al. (2022) calculated seismic masses for several stars and compared the averages of various evolutionary phases to estimate the integrated mass loss between stages in the cluster. M4 also contains several RR Lyrae stars, including overtone ones that have been observed by Kepler. If non-radial modes can be detected in the overtone stars, their seismic masses could be compared directly to the seismic masses of rHB stars, validating the results.
@@ 2.2 Observational constraints
A detailed asteroseismic analysis requires further constraints on the physical parameters (such as L, , or [Fe/H]) of the stars. Howell et al. (2022) used scaling relations for the red giants in the cluster, which relies on luminosities and effective temperatures. These quantities can be calculated from photometry, but they require accurate corrections for interstellar extinction. Here we used the reddening map and E(B–V) zero point determined specifically for M4 by Hendricks et al. (2012).
@@
We collected the average brightness values of the RRc stars both in Gaia DR3 passbands, and in Johnson passbands from the observations of Stetson et al. (2014)222Data available at https://www.cadc.hia.nrc.gc.ca/en/community/STETSON/homogeneous/ and at https://www.canfar.net/storage/list/STETSON/homogeneous/.. The Johnson data set is displayed in Fig. 2, with the RRc stars and the stars studied by Howell et al. (2022) highlighted. However, since input parameters for previous RRc model fits were based on Gaia brightnesses, we decided to use those here, as well.
First, we calculated the interstellar extinction in the Gaia passbands with the seismolab software package, which uses the mwdust code and the Bayestar 3D dust map for that purpose (Bódi et al., 2022; Bovy et al., 2016; Green et al., 2019). We then used the distance modulus provided by Baumgardt & Vasiliev (2021), mag, to convert the extinction-corrected values into absolute magnitudes.
We used and to derive luminosity and effective temperature. The effective temperature was calculated using the relation from eq. 1 from Mucciarelli & Bellazzini (2020), which requires metallicity, [Fe/H], and colour. We employed the relation derived for giant stars and colour . We set the metallicity to [Fe/H] (Harris, 1996; Marino et al., 2008; MacLean et al., 2018). The luminosity was calculated using the formula:
where is a bolometric brightness of a star, and is a bolometric brightness of the Sun, set to 4.74 mag. We converted MG to using bolometric correction calculated with the Gaiadr3 Bcg tool by Creevey et al. (2023). The input parameters to calculate bolometric correction are gravity (log), metallicity, and effective temperature. We assumed log.
@@
where [/Fe] is enhancement by elements. Then, we transformed [M/H] to , using the relation:
@@ 3.3 Comparison with other photometric results
V6 | V42 | V43 | V61 | V76 | |
@@ 3.4 RRc pulsation models
In order to find the best matches between periods and period ratios in theoretical models and in observed values for each star, we employed genetic algorithms333We used a Python geneticlagorithm library https://github.com/rmsolgi/geneticalgorithm, which used the pulsation code with different input physical parameters. The ranges of physical parameters were set based on observational constraints (see Table 1). In the case of mass, we set a wide range of M⊙ to accommodate any realistic mass result. For each star, we executed 100 separate runs. For each run we set the maximum number of iterations to 150, population size of 100, mutation probability of 0.1, elitist ratio to 0.01, crossover probability to 0.5, and parents portion to 0.3. We note that we performed calculations with different parameters for genetic algorithms beforehand, including the numbers of populations and iterations, to ensure that the results are robust. The values and errors of physical parameters were derived as the means and standard deviations of the results from each of the hundred individual runs.
Calculated ranges of physical parameters used to constrain the models for each star are summarized in Table 1.
V06 | V42 | V43 | V61 | V76 | |
M | |||||
X | |||||
Z |
@@ 4.1 RR Lyrae masses and other physical parameters
As listed in Table 2, the masses of four out of the five RRc stars are within 0.06 , between 0.631 and 0.683 with an average uncertainty of . The only outlier is the modulated star, V61, at 0.751 , but for that star our mode identification was more uncertain. However, mass outliers exist among the other evolutionary stages, such as among the rHB stars in the cluster as well. These masses align well with the mass-period distribution of RRc stars published by Netzel et al. (2023).
We compared the calculated and values to the distribution of RR Lyrae stars presented by Molnár et al. (2023b). As Fig. 5 shows, the stars belong to the cooler RRc stars, lying between 6500–7000 K, with values being very close to 3.0 for all five stars. For chemical composition, we find average bulk abundances of and . These values indicate a He abundance of for M4.
@@
4.2 Seismic masses for red giants and the average rHB mass

### v3

@@
11email: molnar.laszlo@csfk.org; lmolnar@konkoly.hu 22institutetext: ELTE Eötvös Loránd University, Institute of Physics and Astronomy, 1117, Pázmány Péter sétány 1/A, Budapest, Hungary 33institutetext: Institute of Physics, École Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, 1290 Versoix, Switzerland 44institutetext: Nicolaus Copernicus Astronomical Centre, Polish Academy of Sciences, Bartycka 18, PL-00-716 Warszawa, Poland 55institutetext: Department of Astronomy, The Ohio State University, 140 W. 18th Ave., Columbus, OH 43210, USA 66institutetext: Center for Cosmology and Astroparticle Physics (CCAPP), The Ohio State University, 191 W. Woodruff Avenue, Columbus, OH 43210, USA 77institutetext: University of Wyoming, 1000 E University Ave, Laramie, WY USA
@@
Globular clusters offer a powerful way to test the properties of stellar populations and the late stages of low-mass stellar evolution. In this paper we study oscillating giant stars and overtone RR Lyrae-type pulsators in the nearest globular cluster, M4, with the help of high-precision, continuous light curves collected by the Kepler space telescope in the K2 mission. We determine the frequency composition of five RRc stars and model their physical parameters from linear pulsation models. We are able, for the first time, to compare seismic masses of RR Lyrae stars directly to the masses of the very similar red horizontal branch stars in the same stellar population, independently determined from asteroseismic scaling relations. We find average seismic masses of for RR Lyrae stars and for red horizontal-branch stars. While the accuracy of our RR Lyrae masses still relies on the accuracy of evolutionary mass differences of neighboring horizontal branch subgroups, this result strongly indicates that RRc stars may indeed exhibit high-degree, and 9 non-radial modes, and modeling these modes can provide realistic mass estimates. We compare the seismic masses of our red horizontal branch and RR Lyrae stars to evolutionary models and to theoretical mass relations and highlight the limitations of these relations.
@@
Binarity, for example, would be a straightforward way to infer dynamical masses. However, not only do we not know any eclipsing binaries with an RR Lyrae member, it is very difficult to find systems that contain an RR Lyrae star at all (Wade et al., 1999; Kennedy et al., 2014; Hajdu et al., 2015; Liška et al., 2016b). Since bona fide RR Lyrae stars are considered to be very old, their companions have either evolved into white dwarfs or other stellar remnants already, or must be smaller stars, most likely still evolving along the lower main sequence as K–M dwarfs. As such, the contribution of RR Lyr companions to the total light output of the system is expected to be minimal, precluding the detection via spectral energy distribution, as double-lined spectroscopic binaries.
The dearth of eclipsing RR Lyrae stars can also be explained by the evolutionary pathways of these pulsators. These stars moved past the red giant phase without any significant interactions with a companion that could have directed the primary away from evolving onto the HB and towards binary evolution scenarios. Companions thus must be distant to avoid any mass transfer, and a distant companion orbiting a now shrunken horizontal-branch star would make the probability of observing the expected shallow eclipses from the system very low. The only eclipsing binary candidate system turned out to be an impostor, a Binary Evolution Pulsator (Pietrzyński et al., 2012; Karczmarek et al., 2017). In this case, mass transfer directed a much lighter primary to rapidly cross the instability strip and to experience RR Lyrae-like pulsations. We note that binarity and a limited amount of mass transfer has been proposed to explain metal-rich RR Lyrae stars, but those systems are expected to have wide orbits, too (Bobrick et al., 2024; Abdollahi et al., 2025).
More indirect methods have shown increased promise, but all have their own caveats. For example, radial velocity searches are hindered by the expected long orbital periods and small signals relative to the much larger pulsation velocities, and TU UMa remains the only viable candidate (Liška et al., 2016a). A recent work proposed new binary candidates, but those still require follow-up and confirmation (Barnes et al., 2021).
An alternative way to detect binaries and estimate masses is the light-time effect, but this method has been impeded by the prevalence of the Tseraskaya–Blazhko effect111Ms. Tseraskaya (Ceraski) collected the data at Moscow Observatory that led to the discovery., the quasi-periodic modulation of the pulsation amplitude and phase (Blažko, 1907; Shapley, 1916; Kurtz, 2022). Nevertheless, many Galactic stars show cyclic phase variations that can be caused by light-time effect (see, e.g., Liška et al., 2016b; Sylla et al., 2024). Furthermore, several binary candidates have been identified in the OGLE survey this way (Hajdu et al., 2015, 2021). And while Hajdu et al. (2021) did not address the masses of the primaries directly, these authors found a trimodal distribution in the mass functions of the secondaries.
@@
One way to limit uncertainties in stellar parameters is to study RR Lyrae stars in globular clusters. There are currently close to three thousand RR Lyrae stars known in 115 Galactic globular clusters (Cruz Reyes et al., 2024). For these stars, parameters like distance and metallicity can be determined accurately for the members, and using that information, the van Albada & Baker (1971) relation was employed in many cases to estimate masses. M3, for example, was studied by multiple authors, and Valcarce & Catelan (2008) determined the mass distribution for the entire HB of the cluster based on evolutionary models, and found the distribution to be potentially bimodal. However, they discuss how that finding can be complicated by limitations present in the stellar tracks. A different approach was taken by Kumar et al. (2024), who fitted the light curves of the M3 RR Lyrae variables directly, based on a grid of non-linear pulsation models. Direct light curve fitting is a promising technique (see also, Bellinger et al., 2020; Das et al., 2025), but current one-dimensional non-linear pulsation models still cannot reproduce light curves very accurately, due to limitations in handling convection, for example (Kovács et al., 2023).
Uncertainties in evolutionary and non-linear pulsation models can also be mitigated if we turn towards asteroseismology, and model the frequencies of the oscillatory modes observed in the star (see, e.g., Kjeldsen & Bedding, 1995; Aerts, 2021). Incorporating asteroseismic constraints has shown great potential recently for stars outside the instability strip. However, classical pulsators like RR Lyrae stars face obstacles in this aspect, too: in these stars, most often only a single mode can be observed, severely limiting our ability to constrain physical parameters from frequencies alone. Masses for a few double-mode RR Lyrae stars have been calculated from non-linear non-adiabatic hydrodynamical simulations, (Molnár et al., 2015b), but, as mentioned before, the accuracy of non-linear models have been called into question (Smolec & Moskalik, 2008).
More recently, a new method based on the discovery of low-amplitude signals in overtone RR Lyrae stars has been proposed (Netzel et al., 2015b, c). Theoretical predictions by Dziembowski (2016) suggested that these modes may be high-order, , 9 non-radial modes, which can be calculated from linear pulsation models. Unlike modeling radial modes in the non-linear regime, fitting radial and non-radial modes together using linear models is a technique very similar to the asteroseismology of red giant stars. This method was developed by Netzel & Smolec (2022), and was further explored, incorporating other observables such as , L and [Fe/H] information into the fit in Netzel et al. (2023).
@@
This opportunity came with the Kepler space telescope. Kepler observed multiple globular clusters during the K2 mission, including the cluster closest to the Sun, Messier 4 (M4). This cluster was bright enough that seismic data was obtainable not only for the bright red giants, but fainter stars, such as the rHB stars in it, as well (Miglio et al., 2016; Tailo et al., 2022). Howell et al. (2022) calculated seismic masses for several stars and compared the averages of various evolutionary phases to estimate the integrated mass loss between stages in the cluster. M4 also contains several RR Lyrae stars, including overtone ones that have been observed by Kepler. If non-radial modes can be detected in the overtone stars, their seismic masses could be compared directly to the seismic masses of rHB stars, validating the results.
@@ 2.2 Observational constraints
A detailed asteroseismic analysis requires further constraints on the physical parameters (such as L, , or [Fe/H]) of the stars. Howell et al. (2022) used scaling relations for the red giants in the cluster, which relies on luminosities and effective temperatures. These quantities can be calculated from photometry, but they require accurate corrections for interstellar extinction. Here we used the reddening map of Alonso-Garc´ıa et al. (2012), but with the reddening zero point (E(B–V)0 = 0.37 mag) determined specifically for M4 by Hendricks et al. (2012).
@@
We collected the average brightness values of the RRc stars both in Gaia DR3 passbands, and in the Johnson B and V passbands from the observations of Stetson et al. (2014, 2019)222Data available at https://www.cadc.hia.nrc.gc.ca/en/community/STETSON/homogeneous/ and at https://www.canfar.net/storage/list/STETSON/homogeneous/.. The Johnson data set is displayed in Fig. 2, with the RRc stars and the stars studied by Howell et al. (2022) highlighted. Although former studies by Netzel & Smolec (2022); Netzel et al. (2023) used Gaia data, here we decided to use on the Johnson brightness for consistency with the red giant analysis.
We calculated the insterstellar extinction for the RRc targets similarly to that of the red giant targets, initially relying on the Alonso-Garc´ıa et al. (2012) and Pancino et al. (2024) differential reddening maps and the same zero point. The latter covers a larger area, but otherwise we found the two maps to be consistent with each other, with differences not exceeding mag. We determined extinctions for four out of five RRc stars this way. We used the zero point of Hendricks et al. (2012) here, too.
The fifth star, V43, falls outside each map. For this star, we compared the E(B–V) reddening values to the E(BP–RP) reddenings from Gaia DR3. We limited the comparison to sources with mag, (BP–RP) ¡ 1.4 mag, and E(BP–RP) ¡ 1.3 mag, and cut out the core of the cluster. These filters were necessary to remove sources with very high reddenings in the catalog. This local comparison of 67 stars resulted in the conversion of . We note, however, that this formula has not been tested more extensively, and such conversions elsewhere will require further calibrations. Furthermore, the RR Lyrae variables themselves have Gaia reddenings considerably larger than their neighboring stars, so we instead calculated the median of nine neighboring stars for V43.
We then used the distance modulus provided by Baumgardt & Vasiliev (2021), mag, to convert the extinction-corrected values into absolute magnitudes.
The effective temperature, , was calculated using the color- relation by González Hernández & Bonifacio (2009) with (V-K)0 colors. For four stars, we used the color indices using -band photometry from 2MASS (Skrutskie et al., 2006). V43 lacks -band photometry: for that star we set the range of to dex. The dereddened color was calculated using the color excess obtained from with the transformation relation of Fitzpatrick & Massa (2007). The calculated ranges of are collected in Table 1.
We set the metallicity to [Fe/H] (Harris, 1996; Marino et al., 2008; MacLean et al., 2018). The luminosity was calculated using the formula:
where is a bolometric brightness of a star, and is a bolometric brightness of the Sun, set to 4.74 mag. We converted MV to using the bolometric correction calculated with the relation by Alonso et al. (1999), which uses and metallicity [Fe/H]. The resulting luminostiy ranges are collected in Table 1.
@@
where [/Fe] is enhancement by elements. Metal-poor populations have elevated element abundances: for M4, Marino et al. (2008) found [/Fe] = (Johnson et al., 2014; Bensby et al., 2017).
Then, we transformed [M/H] to , using the relation:
@@ 3.3 Comparison with other photometric results
We also analyzed the observations collected by the Zwicky Transient Facility (ZTF, Bellm et al., 2019) and the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al., 2018). The ZTF data was too sparse for a detailed frequency analysis. The ATLAS survey collected significantly more data points, but the photometric precision of the observations was too low to detect any extra modes.
| ID | V6 | V42 | V43 | V61 | V76 |
@@ 3.4 RRc pulsation models
In order to find the best matches between periods and period ratios in theoretical models and in observed values for each star, we employed genetic algorithms333We used a Python geneticlagorithm library https://github.com/rmsolgi/geneticalgorithm, which used the pulsation code with different input physical parameters. The ranges of physical parameters were set based on observational constraints (see Table 1). In the case of mass, we set a wide range of M⊙ to accommodate any realistic mass result. For each star, we executed 50 separate runs. For each run we set the maximum number of iterations to 150, population size of 100, mutation probability of 0.1, elitist ratio to 0.01, crossover probability to 0.5, and parents portion to 0.3. We note that we performed calculations with different parameters for genetic algorithms beforehand, including the numbers of populations and iterations, to ensure that the results are robust. The values and errors of physical parameters were derived as the means and standard deviations of the results from each of the fifty individual runs.
Calculated ranges of physical parameters used to constrain the models for each star are summarized in Table 1. Correlations between individual parameters based on the distributions of the results from each run are visualized with corner plots and discussed in Appendix C.
| ID | V06 | V42 | V43 | V61 | V76 |
| M | () | ||||
| X | () | ||||
| Z | () |
@@ 4.1 RR Lyrae masses and other physical parameters
As listed in Table 2, the masses of four out of the five RRc stars are within 0.03 , between 0.636 and 0.667 with an average uncertainty of . These masses align well with the mass-period distribution of RRc stars published by Netzel et al. (2023).
The only outlier is the modulated star, V61, for which we found no solutions with unstable non-radial modes. This is caused by the higher value of the star which puts it near the blue edge of the RR Lyrae instability strip, an outside of the instability region of the 0.61 modes (Netzel & Smolec, 2022). This suggests that the frequency component we detected and identified tentatively as might have a different origin. If we allow for solutions with stable modes at the detected mode frequency, we find a mass of which is significantly higher than the rest. We include this value for the sake of completeness, but consider it as a marginal detection and do not include it in our statistics.
We compared the calculated and values to the distribution of RR Lyrae stars presented by Molnár et al. (2023b). As Fig. 5 shows, the stars belong to the cooler RRc stars, lying between 6600–7200 K, with values being very close to 3.0 for all five stars. For chemical composition, we find average bulk abundances of and . These values indicate a He abundance of for M4, pointing towards low He enrichment () relative to the primordial He abundance, which is in tension with the Y values inferred by Valcarce et al. (2014) spectroscopically. On the other hand, our result is in agreement with the model-based inferences of Villanova et al. (2012), supporting their conclusion that the spectroscopic data should be reevaluated, and followed upon by further observations. We note, however, that the distribution of X values among the individual model runs are skewed towards higher values for two stars (V42, V76), but are bimodal for the other two (V06 and V43) as shown in Appendix C. For V61 the fits skew towards lower values but since it is a marginal detection we do not consider it in the averages.
@@
4.2 Seismic masses for red giants